# Boosting Adversarial Transferability for Hyperspectral Image Classification Using 3D Structure-invariant Transformation and Intermediate Feature Distance

### v1

@@ IV Experiments
In this section, we detail the implementation details and experimental results of our method on two public datasets. First, we introduce the dataset information, parameter settings, and comparison methods. Next, we evaluate the transferability of the adversarial examples on black-box models and their robustness against two defense methods. Finally, we analyze the impact of several hyperparameters on the experimental results.
@@ IV-A Experimental Setup
1) Datasets: Experiments were conducted on two widely recognized public datasets: HoustonU 2018 and PaviaU, as shown in Fig. 4.
Refer to captionRefer to captionRefer to caption
Figure 4: False color images and ground-truth maps of the (a) HoustonU 2018 and (b) PaviaU datasets.
TABLE II: Training and testing sets of the HoustonU 2018
and PaviaU datasets
HoustonU 2018 PaviaU
Class Train Test Class Train Test Class Train Test
1 5000 4799 11 1000 24002 1 900 5731
2 10000 22502 12 100 516 2 900 17749
3 500 184 13 10000 36358 3 900 1199
4 10000 3588 14 5000 4849 4 900 164
5 3000 2048 15 5000 1937 5 900 445
6 3000 1516 16 10000 1475 6 900 4129
7 200 66 17 100 49 7 900 430
8 10000 29762 18 3000 3578 8 900 2782
9 10000 213684 19 3000 2365 9 900 47
10 10000 35810 20 3000 3824
HoustonU 2018: This dataset was captured by the HSI CASI 1500 sensor in the Houston University area on February 16, 2017. It has a spatial resolution of 1 meter, a spatial size of 601 × 2384 pixels, 48 spectral bands, and 20 available classes of interest. The dataset can be obtained by applying through the official website.
PaviaU: This dataset was collected by the ROSIS-03 satellite over the Pavia University area. It has a spatial resolution of 1.3 meters, a spatial size of 610 × 304 pixels, and includes 115 spectral bands. Prior to the experiments, 12 noisy spectral bands were removed, leaving 103 usable spectral bands and 9 available classes for classification.
2) Training examples: For each pixel, a neighborhood window with a spatial size of 32 × 32 is applied to construct the HSIs examples. The detailed number of training and testing examples for both datasets is provided in Table II.
3) Networks: We use five pretrained classification networks to verify the transferability of the attack: VGG-11, Resnet-18, VGG-19, Inception v3 (Inc-v3), and Inception Resnet V2 (IncRes-v2). Consistent with the paper [33], VGG-11 and Resnet-18 served as white-box models to generate adversarial examples, and other models are used to test transferability. In addition, we also choose two common defense methods based on image processing, which are random noise addition and spectral smoothing filtering, to test the robustness of the attack.
4) Attack Methods: We use FGSM and MI-FGSM as baseline attack methods, and combine our method with baseline attack methods, namely ours-FGSM and ours-MI-FGSM. Additionally, the study [33] addresses related issues and is denoted as SS-FGSM and SS-MI-FGSM. The most relevant work [35] is denoted as SIA-FGSM and SIA-MI-FGSM, respectively. Our method is compared against these approaches to demonstrate its effectiveness.
TABLE III: Intermediate layers used for feature extraction. ”Block” refers to the basic blocks in ResNet-18. ”ReLU-k” denotes the k-th ReLU layer in VGG-11
@@
5) Parameters: In this paper, two smaller disturbance budgets are selected for experiments, which are 0.01 and 0.03 respectively. Some parameters of the SS-FGSM method follow the settings of the original paper: the spatial superpixel size is set to four spatial pixels, and the spectral clustering size is set to 20. The ground sample distances (GSDs) of the HoustonU 2018 dataset is 4 m, and that of the PaviaU dataset is 5.2 m. The number of external iterations $T=20$, $\mu=1$. Considering that the improvement in accuracy is not evident relative to the computational complexity, the parameter $t$ is set to 1. To ensure fairness with other methods, the main parameters of our method are established as follows: the spatial division parameter is set to 3, consistent with SIA, and the spectral division parameter is also set to 3. The iteration count is set to 20, the number of transformed example copies is 10 and the feature loss weight coefficient $n$ is set to 0.1. For the white-box model, one layer is selected for feature extraction from shallow to deep layers (1/2/3/4, as shown in Table III). In our experiments, both models selected layer 3. The original feature enlargement factor is set to 1.2, the momentum factor for MI is set to 1, the step size is 2/255, and the batch size is 100. The experiments are carried out with 13th Gen Intel Core i5-13600KF CPUs with 32 GB of RAM, and one NVIDIA GeForce RTX 4090 GPU.
@@
IV-B Transferability Experiments
In this section, we evaluate the performance of the proposed method on five pretrained classification models. The evaluation metrics include Overall Accuracy (OA), Average Accuracy (AA), and the Kappa coefficient. Lower values of these metrics indicate a lower recognition accuracy of the model on adversarial examples, thereby reflecting better adversarial attack performance. The experimental results on the HoustonU 2018 dataset are presented in Tables IV and V, while the results on the PaviaU dataset are shown in Tables VI and VII. The maximum perturbation budgets used were 0.01 and 0.03. The experimental results represent the averages of multiple trials to reduce the influence of randomness introduced by random operations.
HoustonU 2018 Dataset: The experimental results are shown in Tables IV and V. Adversarial examples were generated using the white-box models ResNet-18 and VGG-11, respectively. It can be seen that across all methods and models, the attack effect of the perturbation budget of 0.03 is significantly better than the perturbation budget of 0.01, indicating that smaller perturbation budget pose greater challenges. When combined with MI-FGSM, our method achieved comparable or even better improvements at a budget of 0.01 compared to 0.03. This indicates that our method can effectively enhance attack performance even under smaller perturbation budgets. The attack performance of our method across all models is significantly better than that of the baseline method. For a perturbation budget of 0.01, the average improvement was 14.06%, while for a budget of 0.03, the average improvement was 25.5%. Compared to the SIA method on four black-box models, the attack success rate increased by an average of 2.44% at $\delta=0.01$ and 3.28% at $\delta=0.03$. However, we observed that the white-box attack success rate decreased significantly compared to SIA. Data from ablation experiments indicate that this is caused by spectral-wise block transformations and feature distancing loss. We believe that the possible reason is that the use of multi-dimensional block transformations and weighted Chi-squared distance loss alter the adversarial example optimization process. This approach not only focuses on the classification boundary of the surrogate model but also pay more attention to the change of feature distribution. Consequently, this may lead to a more dispersed disturbance in the input space, thereby diminishing the effectiveness of attacks on the substitute model and resulting in a poorer performance in white-box attacks.
@@ IV-B Transferability Experiments
PaviaU Dataset: The experimental results are shown in Tables VI and VII. We present the results of adversarial examples generated on the ResNet-18 and VGG-11 models. Unlike the HoustonU 2018 dataset, attacking black-box models on this dataset is more challenging due to its higher number of spectral bands. On this dataset, our method largely retained the attack success rate on white-box models while significantly improving baseline attack transferability. Moreover, the performance improvement on the PaviaU dataset was slightly better than that on the HoustonU 2018 dataset. Compared to the baseline method across all models, the attack success rate of our method increased by an average of 13.12% at $\delta=0.01$ and 18.07% at $\delta=0.03$. Compared to the SIA method across four black-box models, the attack success rate increased by an average of 2.05% at $\delta=0.01$ and 7.31% at $\delta=0.03$. Compared to ResNet-18, the adversarial examples generated when VGG-11 is used as a substitute model show greater improvements in attack performance. We believe that this is due to differences in model structure, feature extraction mechanisms, and attention areas, which lead to varying attack effects.
IV-C Defense Experiments
To verify the robustness of the proposed method against common defense strategies, we conducted experiments on two public datasets using two defense methods: adding random noise and spectral dimension filtering. Both defense techniques are commonly used in image processing to mitigate or disrupt adversarial perturbations, thereby improving model robustness. Comparisons with the baseline, SS, and SIA methods are presented in Tables VIII and IX.
Adding random noise: Before feeding examples into the model, random noise is added to the examples to disrupt the structure of adversarial perturbations, reducing the model’s sensitivity to adversarial examples and thus weakening the effectiveness of adversarial attacks. The budget for the added random noise is set at 0.1, and the results are shown in Table VIII. It can be observed that our proposed method maintains good attack performance even under the random noise defense.
Spectral filtering: HSIs contain redundant spectral bands; therefore, local spectral filtering is also a challenging defense method. We perform average filtering along the spectral dimension to validate the robustness of the attack method. The filter size is set to 7, and the results are presented in Table IX. Our method demonstrates strong adaptability and robustness across different datasets and defense strategies.
@@ IV-C Defense Experiments
TABLE VIII: Overall classification accuracy (%) of HSI adversarial examples under added random noise conditions. The adversarial examples for the two datasets are generated on VGG-11 with a maximum perturbation budget of 0.03, and the intensity of the random noise is set to 0.1
@@
IV-D Ablation experiment
In this section, we explore the influence of each component of the generation method on attack performance through ablation experiments. Additionally, we investigate how various hyperparameters affect the experimental results.
TABLE IX: Overall classification accuracy (%) of HSI adversarial examples under spectral filtering conditions. The adversarial examples for the two datasets are generated on VGG-11 with a maximum perturbation budget of 0.03, and the filter size range is set to 7
@@ IV-D Ablation experiment
First, we analyze the impact of each component on the attack effectiveness. The experimental results are summarized in Table X. The relevant parameters have been described in the previous section on parameter settings. Although performance slightly varies across different models, the attack effectiveness generally improves with the application of our method. The results show that, compared to the spatial dimension, performing block transformations in the spectral dimension can also reduce the model’s classification accuracy. Combining both approaches improved the transferability on the two datasets by approximately 4% and 6%, respectively. Additionally, integrating feature distancing loss further enhanced the attack transfer success rate of adversarial examples. The performance on the PaviaU dataset was better than on the HoustonU 2018 dataset, which may be related to the data distribution and the size of the regions on which the model focuses. Finally, using the enlarged original example features when calculating the Chi-squared distance can further improve the attack success rate, thereby validating the effectiveness of the method.
TABLE X: Ablation study of the proposed approach. The adversarial examples are crafted on VGG-11, with the integrated baseline method being FGSM. The perturbation budget is set to 0.03. SpatialT represents spatial transformation; SpectralT represents spectral transformation; FDL represents feature distancing loss; and Enlarge represents the operation of enlarging the features of the original sample

### v2

@@ IV Experiments
In this section, we detail the implementation details and experimental results of our method on three public datasets. First, we introduce the dataset information, parameter settings, and comparison methods. Second, we evaluate the transferability of the adversarial examples on black-box models and their robustness against two defense methods. Subsequently, we discuss the impact of multiple parameters on experimental results. Finally, we analyze a set of visualization results.
@@ IV-A Experimental Setup
1) Datasets: Experiments were conducted on three widely used public datasets: HoustonU 2018, PaviaU, and Indian Pines as shown in Fig. 5.
HoustonU 2018: This dataset was captured by HSI CASI 1500 sensor in the Houston University area on February 16, 2017. It has a spatial resolution of 1 meter, a spatial size of 601 × 2384 pixels, 48 spectral bands, and 20 available classes of interest. The dataset can be obtained by applying through the official website.
PaviaU: This dataset was collected by ROSIS-03 satellite over the Pavia University area. It has a spatial resolution of 1.3 meters, a spatial size of 610 × 304 pixels, and includes 115 spectral bands. Prior to the experiments, 12 noisy spectral bands were removed, leaving 103 usable spectral bands and 9 available classes for classification.
Indian Pines: This dataset was collected over the Indianapolis area of Indiana, USA, using the AVIRIS sensor. It has a spatial resolution of 20 meters, a spatial size of 145 × 145 pixels, and includes 200 spectral bands, and 16 available classes of interest.
2) Training examples: For each pixel, a neighborhood window with a spatial size of 32 × 32 is applied to construct the HSIs examples. The detailed number of training and testing examples for the three datasets are provided in Table II.
3) Networks: We use five pretrained classification models to verify the transferability of the attack: VGG-11, ResNet-18, VGG-19, Inception v3 (Inc-v3), and Inception Resnet V2 (IncRes-v2). Consistent with the work of SS [32], VGG-11 and ResNet-18 served as substitute models to generate adversarial examples, and other models are used to test transferability.
4) Attack Methods: We use FGSM [18] and MI-FGSM [20] as the baseline attack methods, and combine our method with baseline attack methods, namely ours-FGSM and ours-MI-FGSM. Other advanced comparison methods are SS [32], SIA [58], and BSR [40], which are also combined with the baseline methods. We compare our method with these approaches to demonstrate its effectiveness. In addition, we also choose two common defense methods based on image processing, which are random noise addition and spectral smoothing filtering, to test the robustness of the attack.
TABLE III: Intermediate layers used for feature maps extraction. ”Block” refers to the basic blocks in ResNet-18. ”ReLU-k” denotes the $k$-th ReLU layer in VGG-11
@@
5) Parameters: In our work, two smaller disturbance budgets are selected for experiments, which are 0.01 and 0.03 respectively. Some parameters of the SS [32] method follow the settings of the original paper: the spatial superpixel size is set to four spatial pixels, and the spectral clustering size is set to 20. The number of external iterations TT=20, $\mu$=1, the parameter $t$ is set to 1. To ensure fairness with other methods, the main parameters of our method are established as follows: the spatial division parameter is set to 3, consistent with SIA [58], and the spectral division parameter is also set to 3. The iteration steps is set to 20, the number of transformed example copies is 10 and the cross-entropy classification loss weight coefficient $\eta$ is set to 0.03. For the substitute model, one layer is selected for feature maps extraction from shallow to deep layers (1/2/3/4, as shown in Table III). In our experiments, both models selected layer 3. The original feature enlargement factor $\lambda$ is set to 1.2, the momentum factor for MI is set to 1, the step size is 2/255, and the batch size is 100. The experiments are carried out with 13th Gen Intel Core i5-13600KF CPUs with 32 GB of RAM, and one NVIDIA GeForce RTX 4090 GPU.
@@
IV-B Transferability Experiments
In this section, we evaluate the performance of the proposed method on five pretrained classification models. The evaluation metrics include Overall Accuracy (OA), Average Accuracy (AA), and the Kappa coefficient. Lower values of these metrics indicate a lower recognition accuracy of the model on adversarial examples, thereby reflecting better adversarial attack performance. In the following experimental analysis, we use the decrease in overall accuracy to reflect the increase in attack success rate. The experimental results on the HoustonU 2018 dataset are presented in Tables IV and V. Results on the PaviaU dataset are shown in Tables VI and VII, while those from the Indian Pines dataset are shown in Tables VIII and IX. The experimental results represent the averages of multiple trials to reduce the influence of randomness introduced by random operations.
TABLE VIII: Classification accuracy (%) of adversarial attacks on five trained networks on the Indian Pines dataset. The adversarial examples are crafted on ResNet-18. The maximum perturbation budget is set to 0.01 and 0.03, respectively
Attack ResNet-18 VGG-11 VGG-19 Inc-V3 IncRes-V2
Method Perturbation 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03
budget
NONE OA 99.92 99.92 99.92 99.92 99.75 99.75 99.94 99.94 99.73 99.73
AA 99.95 99.95 99.70 99.70 99.50 99.50 99.93 99.93 99.76 99.76
Kappa 99.92 99.92 99.92 99.92 99.71 99.71 99.93 99.93 99.69 99.69
FGSM[18] OA 91.21 15.10 99.88 93.89 99.59 94.74 99.94 94.57 99.50 92.46
AA 89.41 17.85 99.19 92.30 99.15 95.61 99.93 97.70 99.72 92.89
Kappa 89.98 4.72 99.87 93.01 99.54 94.02 99.93 93.83 99.43 91.39
$S$-FGSM[32] OA 88.91 3.90 99.88 86.71 99.69 91.54 99.94 90.07 99.54 90.99
AA 88.32 6.53 99.19 88.97 99.22 92.40 99.93 94.65 99.73 87.82
Kappa 87.36 -8.70 99.87 85.81 99.65 90.26 99.93 88.61 99.47 89.64
SIA-FGSM[58] OA 99.87 5.14 90.51 91.34 99.58 92.85 99.94 90.79 99.59 92.86
AA 98.57 8.21 90.79 91.11 98.66 93.51 99.93 95.94 99.55 92.26
Kappa 99.85 -4.60 89.21 90.11 99.52 92.19 99.93 89.54 99.54 91.83
BSR-FGSM[40] OA 98.18 23.13 99.90 88.53 99.71 90.52 99.94 85.11 99.59 89.78
AA 98.97 35.05 99.07 93.25 99.45 94.85 99.93 94.03 99.70 91.02
Kappa 97.93 12.72 99.89 86.85 99.67 89.22 99.93 83.14 99.54 88.29
ours-FGSM OA 96.46 18.01 99.90 86.20 99.65 88.21 99.92 82.73 99.63 89.04
AA 94.95 20.05 99.22 91.85 99.21 91.35 99.92 93.85 99.74 91.29
Kappa 95.97 10.86 99.89 84.53 99.60 86.63 99.91 80.63 99.58 87.73
MI-FGSM[18] OA 85.45 4.79 99.88 91.07 99.58 92.10 99.94 90.20 99.50 90.70
AA 85.47 7.23 99.19 92.19 99.15 95.03 99.93 95.50 99.70 90.54
Kappa 83.43 -6.36 99.87 89.75 99.52 90.99 99.93 88.85 99.43 89.35
$S$-MI-FGSM[32] OA 89.08 10.69 99.88 87.86 99.65 92.04 99.94 92.71 99.58 92.93
AA 88.44 15.29 99.19 91.06 99.18 95.40 99.93 95.94 99.77 93.72
Kappa 87.56 -0.22 99.87 86.00 99.60 90.88 99.93 91.64 99.52 91.87
SIA-MI-FGSM[58] OA 89.81 6.51 99.85 90.87 99.56 92.25 99.94 88.93 99.61 92.26
AA 89.83 10.05 98.55 90.83 98.64 93.84 99.93 95.64 99.79 91.48
Kappa 88.43 -3.08 99.82 89.62 99.49 91.20 99.93 87.50 99.56 91.17
BSR-MI-FGSM[40] OA 97.80 24.32 99.88 87.08 99.69 88.63 99.94 80.88 99.58 87.97
AA 97.93 37.06 98.59 93.46 99.42 94.69 99.93 93.46 99.68 93.25
Kappa 97.49 14.58 99.87 85.33 99.65 87.16 99.93 78.59 99.52 86.30
ours-MI-FGSM OA 96.56 19.34 99.88 85.44 99.61 87.57 99.92 83.08 99.67 87.19
AA 94.92 21.57 98.59 90.11 99.18 92.49 99.92 94.21 99.77 90.04
Kappa 96.08 12.32 99.87 83.67 99.56 86.29 99.91 81.07 99.63 85.63
TABLE IX: Classification accuracy (%) of adversarial attacks on five trained networks on the Indian Pines dataset. The adversarial examples are crafted on VGG-11. The maximum perturbation budget is set to 0.01 and 0.03, respectively
Attack VGG-11 ResNet-18 VGG-19 Inc-V3 IncRes-V2
Method Perturbation 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03
budget
NONE OA 99.92 99.92 99.92 99.92 99.75 99.75 99.94 99.94 99.73 99.73
AA 99.70 99.70 99.95 99.95 99.50 99.50 99.93 99.93 99.76 99.76
Kappa 99.92 99.92 99.92 99.92 99.71 99.71 99.93 99.93 99.69 99.69
FGSM[18] OA 95.81 46.71 99.83 97.16 99.75 99.00 99.94 97.58 99.63 93.64
AA 92.64 51.44 99.17 97.59 99.03 98.45 99.93 99.11 99.71 90.26
Kappa 95.21 40.28 99.80 96.77 99.71 98.85 99.93 97.25 99.58 92.77
$S$-FGSM[32] OA 95.88 25.90 99.85 90.24 99.73 95.79 99.94 97.29 99.71 97.26
AA 92.09 29.46 99.65 92.96 99.01 96.78 99.93 98.18 99.75 94.53
Kappa 95.30 16.52 99.82 88.81 99.69 95.19 99.93 96.90 99.67 96.87
SIA-FGSM[58] OA 97.33 23.62 99.77 84.70 99.63 91.20 99.94 88.45 99.54 87.30
AA 92.71 27.52 99.12 91.01 98.70 93.40 99.93 95.01 99.66 91.22
Kappa 96.95 14.65 99.74 82.53 99.58 89.98 99.93 86.86 99.47 85.53
BSR-FGSM[40] OA 99.85 83.70 99.88 94.17 99.73 98.48 99.94 98.21 99.69 91.13
AA 98.77 79.41 99.91 97.41 99.47 98.88 99.93 99.32 99.74 95.94
Kappa 99.82 81.67 99.87 93.40 99.69 98.26 99.93 97.96 99.65 89.97
ours-FGSM OA 98.49 28.65 99.79 69.73 99.63 86.10 99.94 77.44 99.54 83.34
AA 94.76 33.78 99.16 78.78 98.68 89.98 99.93 90.82 99.64 83.95
Kappa 98.28 19.10 99.76 65.53 99.58 84.18 99.93 74.50 99.47 80.93
MI-FGSM[18] OA 78.04 15.68 99.79 92.83 99.71 95.28 99.94 94.86 99.56 90.99
AA 72.87 10.32 99.15 94.50 98.77 95.50 99.93 97.41 99.67 88.82
Kappa 74.91 5.63 99.76 91.81 99.67 94.63 99.93 94.13 99.49 89.76
$S$-MI-FGSM[32] OA 95.54 46.04 99.83 77.17 99.75 92.50 99.94 91.96 99.63 97.68
AA 91.86 54.03 99.17 87.76 99.03 96.12 99.93 96.78 99.71 92.93
Kappa 94.90 39.40 99.80 74.36 99.71 91.46 99.93 90.86 99.58 97.35
SIA-MI-FGSM[58] OA 94.07 14.08 99.67 80.79 99.58 89.07 99.92 83.36 99.46 82.22
AA 89.22 16.00 99.07 89.91 98.64 91.81 99.93 92.75 99.13 85.58
Kappa 93.21 4.48 99.63 78.26 99.52 87.58 99.91 81.20 99.38 79.79
BSR-MI-FGSM[40] OA 99.75 58.68 99.87 88.24 99.69 94.39 99.94 92.34 99.61 80.32
AA 98.22 62.38 99.89 95.03 99.44 97.38 99.93 97.53 99.69 89.49
Kappa 99.71 54.61 99.85 86.81 99.65 93.64 99.93 91.36 99.56 78.05
ours-MI-FGSM OA 96.79 21.99 99.73 67.49 99.63 83.46 99.92 72.43 99.48 77.20
AA 93.04 24.91 99.13 75.89 98.68 87.86 99.93 88.54 99.62 75.50
Kappa 96.33 11.80 99.69 63.16 99.58 81.22 99.91 68.99 99.40 73.56
HoustonU 2018 Dataset: The experimental results are shown in Tables IV and V. Adversarial examples were generated using the substitute models ResNet-18 and VGG-11, respectively. It can be seen that across all methods and models, the attack effect of the perturbation budget of 0.03 is almost significantly better than the perturbation budget of 0.01, indicating that smaller perturbation budget pose greater challenges. When combined with MI-FGSM, our method achieved comparable or even better improvements at a budget of 0.01 compared to 0.03. This indicates that it can effectively enhance attack performance even under smaller perturbation budgets. The attack performance of our method across all models is significantly better than that of the baseline method. In terms of overall accuracy (OA), when the perturbation budget of 0.01, the average improvement was 14.06%, while for a budget of 0.03, the average improvement was 25.5%. Compared to the SIA method on four black-box models, the attack success rate increased by an average of 2.44% at $\varepsilon$ = 0.01 and 3.28% at $\varepsilon$ = 0.03. However, it can be observed that the white-box attack success rate decreased significantly compared to SIA. Data from ablation experiments indicate that this is caused by spectral-wise block transformations and weighted feature divergence loss. The possible reason is that the use of multi-dimensional block transformations and feature loss alter the adversarial example optimization process. Our method not only focuses on the classification boundary of the substitute model but also pays more attention to the change of feature distribution. Consequently, this may lead to a more dispersed disturbance in the input space, thereby diminishing the effectiveness of attacks on the substitute model and resulting in a poorer performance in white-box attacks.
@@ IV-B Transferability Experiments
PaviaU Dataset: The experimental results are shown in Tables VI and VII. We present the results of adversarial examples generated on the ResNet-18 and VGG-11 models. Unlike the HoustonU 2018 dataset, attacking black-box models on this dataset is more challenging due to its higher number of spectral bands. On this dataset, our method largely retained the attack success rate on white-box models while significantly improving baseline attack transferability. Moreover, the performance improvement on the PaviaU dataset was slightly better than that on the HoustonU 2018 dataset. Compared to the baseline method across all models, the attack success rate of this method increased by an average of 13.12% at $\varepsilon$ = 0.01 and 18.07% at $\varepsilon$ = 0.03. Compared to the SIA method across four black-box models, the attack success rate increased by an average of 2.05% at $\varepsilon$ = 0.01 and 7.31% at $\varepsilon$ = 0.03. Compared to ResNet-18, the adversarial examples generated when VGG-11 is used as a substitute model show greater improvements in attack performance. This is due to differences in model structure and feature maps extraction mechanisms, which lead to varying attack effects.
@@
TABLE X: Overall accuracy (%) of HSI adversarial examples under added random noise conditions. The adversarial examples for the three datasets are generated using the VGG-11 model, with the perturbation budget set to 0.03 and the intensity of the random noise set to 0.1
@@
Indian Pines VGG-11 ResNet-18 VGG-19 Inc-V3 IncRes-V2
FGSM 79.20 96.68 98.36 96.06 75.63
$S$-FGSM 70.18 88.85 95.92 95.52 68.61
SIA-FGSM 40.61 83.42 89.80 84.73 66.27
BSR-FGSM 90.38 94.20 96.79 96.83 78.82
ours-FGSM 39.24 67.55 83.46 71.45 55.22
MI-FGSM 60.96 91.90 95.32 94.09 71.94
$S$-MI-FGSM 67.55 76.03 92.81 89.29 68.46
SIA-MI-FGSM 35.93 78.49 86.05 78.20 62.39
BSR-MI-FGSM 73.68 86.63 90.78 89.83 71.16
ours-MI-FGSM 37.11 65.31 79.26 64.48 51.86
TABLE XI: Overall accuracy (%) of HSI adversarial examples under spectral filtering conditions. The adversarial examples for the three datasets are generated using the VGG-11 model, with the perturbation budget set to 0.03 and the filter size range set to 7
@@
Indian Pines VGG-11 ResNet-18 VGG-19 Inc-V3 IncRes-V2
FGSM 70.27 87.57 98.03 95.25 96.85
$S$-FGSM 52.57 83.05 94.80 93.35 97.26
SIA-FGSM 39.53 78.62 91.46 87.03 88.96
BSR-FGSM 89.80 84.19 96.35 97.53 91.55
ours-FGSM 51.04 67.30 86.97 79.36 84.60
MI-FGSM 34.42 85.99 95.34 92.75 94.82
$S$-MI-FGSM 63.24 74.26 91.77 85.08 94.01
SIA-MI-FGSM 36.18 74.84 88.48 84.60 86.32
BSR-MI-FGSM 78.55 76.77 92.29 91.36 83.40
ours-MI-FGSM 44.55 63.66 83.98 75.71 78.32
Indian Pines Dataset: The experimental results are shown in Tables VIII and IX. First, in contrast to the other two datasets, adversarial examples generated on this dataset with a perturbation budget of $\varepsilon$ = 0.01 exhibit no transferability. We believe that this is because the dataset is crop and soil zoning, with small inter-class differences and highly correlated spectra. Different networks will capture different small spectral segments, resulting in low cosine similarity between input gradient direction of the substitute model and black-box model, making small perturbations insufficient to reliably change the prediction. Consequently, we consider $\varepsilon$ = 0.01 not informative for comparison. At $\varepsilon$ = 0.03, this method improves the black-box attack success rate by approximately 12.1% on average over the baseline methods and by about 8.7% over SIA. Compared with the SIA method, the BSR performs better on ResNet-18 but worse on VGG-11, whereas our method is particularly effective when VGG-11 is used as the substitute model.
IV-C Defense Experiments
To verify the robustness of the proposed method against common defense strategies, we conducted experiments on three public datasets using two defense methods: adding random noise and spectral filtering. Both defense techniques are commonly used in image processing to mitigate or disrupt adversarial perturbations, thereby improving model robustness. Comparison results are presented in Tables X and XI.
Adding random noise: Before feeding examples into the model, random noise is added to the examples to disrupt the structure of adversarial perturbations, reducing the model’s sensitivity to adversarial examples and thus weakening the effectiveness of adversarial attacks. The budget of the added random noise is set to 0.1, and the results are shown in Table X. It can be observed that our proposed method maintains good attack performance even under the random noise defense.
Spectral filtering: HSIs contain redundant spectral bands, therefore, local spectral filtering is a challenging defense method. We perform average filtering along the spectral dimension to validate the robustness of the attack method. The filter size is set to 7, and the results are presented in Table XI. Our method demonstrates strong adaptability and robustness across different datasets and defense strategies.
IV-D Ablation Experiment
In this section, we explore the influence of each component of the generation method on attack performance through ablation experiments. The relevant parameters have been described in the previous section on parameter settings. The experimental results are summarized in Table XII. The components in the table are explained as follows: SpatialT refers to block partitioning applied only to the spatial dimension; SpectralT refers to block partitioning applied only to the spectral dimension; WFDL-unEnlarge refers to the use of weighted feature divergence loss without enlarging the original example feature values; WFDL refers to the use of weighted feature divergence loss with enlarging of the original example feature values.
TABLE XII: Ablation study of the proposed approach. The adversarial examples are crafted on the VGG-11 model, with the integrated baseline method being FGSM. The data represents the overall accuracy (%), with the perturbation budget set to 0.03